OpenStudy (anonymous):

How would you factor 4x^2 + 9? I think the answer has imaginary numbers in it, but how would you get the solution?

OpenStudy (anonymous):

4x^2 = -9 x^2 = -9/4 x = +/- SQRT(-9/4) x = +/- 3/2(i)

OpenStudy (anonymous):

Couldn't you just do the square root of both sides to get 2x= SQRT(-9) ?

OpenStudy (accessdenied):

you would get 2x = +- sqrt(-9)

OpenStudy (anonymous):

Oh yeah the +/-, but besides that, that way works?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Thanks :)

OpenStudy (accessdenied):

I suppose theoretically, we could use difference of squares to factor this, asserting that the expression can be rewritten as "4x^2 - 9(i^2)", which would factor into (2x + 3i)(2x - 3i). In practice, however, we would simply say it is not factorable because we usually look for integers when factoring.

OpenStudy (anonymous):

It says to find all real and imaginary factors. So it would be (2x-3i)(2x+3i)

OpenStudy (accessdenied):

ah, well, that may work for your situation then. :P ive never been asked to find imaginary factors for polynomials in my algebra classes. D: